OpenStudy (anonymous):

the height of a triangle is four times the length of the base. find the base and height of the triangle if the area is 18 square inches.

OpenStudy (destinymasha):

12

OpenStudy (anonymous):

thank you for the direct answer but that isnt what i need i would please appreciate it if someone could help me how to do it step by step

OpenStudy (campbell_st):

well you know from the area formula \[18 = \frac{1}{2} b \times h\] so \[36 = b \times h\] so if b = base then the height is h = 4b now your equation becomes \[36 = b \times 4b... or.... 36 = 4b^2\] you need to solve for b

OpenStudy (anonymous):

so when i solve for b all i have to do is divide 36/4

OpenStudy (destinymasha):

yup

OpenStudy (campbell_st):

thats correct, then take the square root of you answer to find b. in measurement questions like this you only consider the positive answer...

OpenStudy (anonymous):

so it will be 3 right. so i times that by 4 to get the height

OpenStudy (destinymasha):

yuppty yup yup

OpenStudy (anonymous):

the base is 3 and height is 12

OpenStudy (destinymasha):

Correct!

OpenStudy (anonymous):

okay thank you both really helped

OpenStudy (destinymasha):

ur welcome